Find the slope and the y-intercept of the line.
2x-4y=-12

Answer by RAY100(1637)

Many different ways to solve this ,,,but lets just match form, y=mx +b
.
2x -4y = -12
.
-4y = -2x -12
.
y = (1/2)x + (3),,,,,or m=1/2,,,b=3
.